OpenStudy (anonymous):

special product of (2x-5y)(2x-5y)

OpenStudy (gorv):

both term are same so like x*x=x^2

OpenStudy (gorv):

similarly it will be have power 2

OpenStudy (anonymous):

No i am asking like 50 sq .r

OpenStudy (shinalcantara):

if you are asking as to what special product is the given expression, then there are three of those for a quadratic expression namely: 1. difference of squares e.g. \[(a ^{2} - b ^{2}) = (a - b)(a + b)\] 2. square of sum e.g. \[a ^{2} + 2ab + b ^{2} = (a + b)^{2}\] 3. square of difference \[a ^{2} - 2ab + b ^{2} = (a - b)^{2}\] what do you think is the expression given?